# bibliography.bib

### v1

@@ C.3 Homogeneous non-decreasing activation with max pooling
Consider the hypothesis space ${\mathcal{H}}$ defined in 4. Suppose that $P(\cdot)$ is the max pooling operation, and $\sigma(\cdot)$ is a $1$-Lipschitz and non-decreasing positive homogeneous activation function. Fix the training data matrix ${\bm{X}}$. Then with probability at least $1-\delta$ and for all $h\in{\mathcal{H}}$, we have:
@@ Appendix F Frequency Domain Analysis
For the hypothesis space ${\mathcal{H}}$ defined in 4, the average pooling operation, $\sigma(\cdot)$ as a $1$-Lipschitz positive homogeneous activation function, the generalization error is bounded as ${\mathcal{O}}(\frac{b_{x}M_{1}M_{2}}{\sqrt{m}})$.
@@
where we used the positive homogeneity property of the activation function, and we defined:
@@ Appendix G Proofs for Weight Sharing
The following works do not consider the generalization error for equivariant networks. However, the convolutional networks, which are studied in their work, are closely related to our results.
@@
Our work is closely related to [vardi_sample_2022]. We focus on equivariant networks and upper bounds on the sample complexity. The networks considered in this work are slightly more general with per-channel linear and non-linear pooling operations. They also provide a bound in Theorem 7 of their paper for pooling operations $\rho:\mathbb{R}^{|G|}\to\mathbb{R}$ that are 1-Lipschitz with respect to the $\ell_{\infty}$ norm. This class includes average and max pooling operations. Their bound has logarithmic dimension dependence, the artifact of using the covering number-based arguments (see below for more discussions). The bound is also dependent on the spectral norm of the matrix ${\bm{W}}$ instead of the norm of the weight ${\bm{w}}$. It is not clear if the dependence on the spectral norm can be removed in their proof. For local filters, they provide a dimension-free bound that also depends on the spectral norm of ${\bm{W}}$. However, this time, the dependence on the spectral norm can be substituted with the norm of the weight vector, as it is evident from their proof. In general, all our bounds depend only on the norms of the weight vector, which is tighter that using Frobenius or spectral norm of ${\bm{W}}$. The authors in [vardi_sample_2022] provide lower bounds on the sample complexity, which we relegate to future works.
@@
The paper of [graf_measuring_2022] relies mainly on Maurey’s empirical lemma and covering number arguments. Rademacher complexity bounds either work directly on bounding the Rademacher sum through a set of inequalities or utilize chaining arguments and Dudley’s integral inequality. Dudley’s integral requires the covering number estimate. The main techniques of finding the covering number, like volumetric, Sudakov, and Maurey’s lemma techniques, manifest different dimension dependencies, with the latter usually providing the mildest dependence. Completely dimension-free bounds are obtained mainly via the direct analysis of the Rademacher sum. Therefore, the bounds in [graf_measuring_2022] are a mixture of dimension-dependent and norm-dependent terms. Their bounds are, however, quite general and consider different activation functions and residual connections.
@@
In [wang_theoretical_2023], they study the same set of problems as ours, related to locality and weight sharing, but from different perspectives. First, the approximation theory perspective, studying the class of functions a model can approximate, is complementary to our paper on the generalization error using statistical learning theory. The learning theory part of the paper is focused on a particular "separation task" (see (5)) with a fixed input distribution. They also use Rademacher complexity but bound it with a covering number. We work with general input distributions and arbitrary tasks. Besides, our bounds are dimension independent, while their bound has explicit dependence on the input dimension, an artifact of using covering number for bounding RC.
@@
In [li_why_2021], the authors consider the sample complexity of convolutional and fully connected models and constructs a distribution for which there is a fundamental gap between them in the sample complexity. Their analysis is based on the VC dimension and uses Benedek-Itai’s lower bound. Their final bound is, therefore, dependent on the input dimension. In contrast, one of the main interests of our paper was to explore dimension-free bounds. We additionally study the impact of locality and weight sharing.
@@
Dimension-free bounds are interesting because they hint at why the generalization error is not affected by over-parametrizing the model. In this sense, the notion of dimension refers mainly to the input dimension, the number of channels, the width of a layer, and the number of layers. The same motivation existed for obtaining norm-based bounds for RKHS SVM models. The works of [golowich_size-independent_2018] and [vardi_sample_2022] are two recent examples, with the former work providing an extensive review of other dimension-free bounds. The choice of a proper norm to get dimension-free bounds is one of the main questions in learning theory. Naturally, this is easier to get for larger norms like Frobenius norms, and the question is if we can have similar dimension-free results for smaller norms like spectral or, as we show in the paper, the norm of effective parameters.
@@
In the literature around deep learning theory, there are many works around questioning the relevance of classical learning theoretic results, for example using Rademacher complexity, norm-based bounds, and uniform convergence. We can refer for example to works such as [nagarajan_uniform_2019, jiang_fantastic_2020, zhang_understanding_2017, zhu_understanding_2021]. It is fair to ask if Rademacher complexity analysis, in light of this criticisms, can provide tight bounds for our case. First of all, the preceding works focused on state of the art deep neural networks with large number of layers. Our study is limited to single-hidden layer neural networks. Second, Rademacher complexity is indeed tight for support vector machines [Steinwart2008-jo, shalev-shwartz_understanding_2014, Bartlett2002-gt]. If we fix the first layer of our network and only train the last layer, then we have a linear model, and it is known that training it with Gradient descent with lead to minimum $\ell_{2}$-norm solutions. For these sub-class of learning algorithms, the analysis boils to previous works on minimum norm classifiers as in [shalev-shwartz_understanding_2014] for which the Rademacher complexity analysis is tight. Therefore, we believe that Rademacher complexity can still be a relevant tool for shallower models including those discussed in this paper.
@@
Previous works also tried to characterize the effect of equivairance on the training dynamics. [lawrence2022implicit] showed that deep equivariant linear networks trained with gradient descent provably converge to solutions which are sparse in the Fourier domain. While this work focuses on the effect of equivariance on optimization, the authors suggest that this implicit bias towards sparse Fourier solutions can be beneficial towards generalization in bandlimited datasets - which is common for non-discrete compact groups such as rotation groups and a common assumption in steerable CNNs. Similar results were also previously obtained in [gunasekar2018implicit] for standard convolutional networks, which can be seen as an instance of GCNN with discrete abelian groups.
@@
In this paper, we provided a Rademacher complexity based bound for single-hidden layer equivariant networks. We obtained a lower bound showing the tightness of our analysis. The numerical results showed compelling scaling behaviors for our bound. However, there is still a gap between our bound and the generalization error. Our conjecture is that the gap is related to the idea of implicit bias. We conjecture that stochastic gradient descent training implicitly biases toward only a subset of the hypothesis space even smaller than norm-bounded functions. It is an important next step to tie these two analysis together by characterizing the implicit bias, for example in terms of some kind of norms, and then use Rademacher complexity analysis to find a bound on the generalization error based on the implicit bias.
On the other hand, we provided a lower bound on the Rademacher complexity. It would be, however, interesting to obtain such bounds by finding the fat-shattering dimension. If two points in the dataset can be transformed into each other by the action of group $G$, then the space of $G$-invariant functions cannot shatter such datasets. Therefore, the data points should be picked on different orbits, so they can be shattered. Such a construction is an interesting future work.

### v2

@@ C.4 Homogeneous non-decreasing activation with max pooling
Consider the hypothesis space ${\mathcal{H}}$ defined in eq. 4. Suppose that $P(\cdot)$ is the max pooling operation, and $\sigma(\cdot)$ is a $1$-Lipschitz and non-decreasing positively homogeneous activation function. Fix the training data matrix ${\bm{X}}$. Then with probability at least $1-\delta$ and for all $h\in{\mathcal{H}}$, we have:
@@ C.5 General Pooling
C.6 Proof for Multi-Layer Group Convolutional Networks
In this section, we prove Theorem 4.3. We start again by simply peeling off the last layer:
${\mathbb{E}_{\bm{\epsilon}}\left(\sup_{{\bm{u}},\{{\bm{w}}^{(l)},l\in[L]\}}\frac{1}{m}\sum_{i=1}^{m}{\epsilon}_{i}{\bm{u}}^{\top}P\circ{\sigma\left({\bm{W}}^{L}{\bm{x}}_{i}^{(L-1)}\right)}\right)}$ $={\frac{M_{1}}{m}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}P\circ\sigma({\bm{W}}{\bm{x}}^{(L-1)}_{i})\right\|\right),}$ (28)
where, for brevity, we use ${\bm{x}}^{(L-1)}$ to denote the output of the hidden layer $L-1$. We can expand the average pooling operation and use the positive homogeneity property of ReLU function to
${\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}P\circ\sigma({\bm{W}}{\bm{x}}^{(L-1)}_{i})\right\|\right)\leq\sum_{g=1}^{|G|}\frac{1}{|G|}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\begin{pmatrix}\sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(1,k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \vdots\\ \sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(c_{L},k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \end{pmatrix}\right\|\right)}$
and then, we focus on a single term and use simple Jensen to get:
$\mathbb{E}_{\bm{\epsilon}}$ $\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\begin{pmatrix}\sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(1,k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \vdots\\ \sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(c_{L},k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \end{pmatrix}\right\|\right)$
$\leq\sqrt{\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\begin{pmatrix}\sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(1,k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \vdots\\ \sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(c_{L},k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \end{pmatrix}\right\|^{2}\right)},$
and we can continue with a similar approach to simplify this further:
$\mathbb{E}_{\bm{\epsilon}}$ $\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\begin{pmatrix}\sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(1,k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \vdots\\ \sigma\left(\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(c_{L},k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\\ \end{pmatrix}\right\|^{2}\right)$
$=\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L]\}}\sum_{j=1}^{c_{L}}{\left\|{\bm{w}}^{(L)}_{(j,:)}\right\|^{2}}\left(\sum_{i=1}^{m}{\epsilon}_{i}\sigma\left(\frac{1}{\left\|{\bm{w}}^{(L)}_{(j,:)}\right\|}\sum_{k=1}^{c_{L-1}}{\bm{w}}_{(j,k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\right)^{2}\right)$
$\leq M_{L}^{2}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1],\tilde{{\bm{w}}^{(L)}}:\left\|\tilde{{\bm{w}}^{(L)}}\right\|\leq 1\}}\left(\sum_{i=1}^{m}{\epsilon}_{i}\sigma\left(\sum_{k=1}^{c_{L-1}}\tilde{{\bm{w}}}_{(k)}^{(L),\top}\Pi_{g}{\bm{x}}^{(L-1)}_{i}{(k)}\right)\right)^{2}\right)$
$\leq M_{L}^{2}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}\left({\bm{x}}^{(L-1)}_{i}\right)\right\|^{2}}\right),$
where $\Pi_{g}$ is a unitary matrix and was removed. Now, we have removed the last layer from the bound, and we can focus on the rest of layers.
$\mathbb{E}_{\bm{\epsilon}}$ $\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}\left({\bm{x}}^{(L-1)}_{i}\right)\right\|^{2}}\right)$
$=\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\begin{pmatrix}\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(1,k)}^{(L-1),\top}\circledast{\bm{x}}^{(L-2)}_{i}{(k)}\right)\\ \vdots\\ \sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(c_{L-1},k)}^{(L-1),\top}\circledast{\bm{x}}^{(L-2)}_{i}{(k)}\right)\\ \end{pmatrix}\right\|^{2}\right)$
$=\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\sum_{g=1}^{|G|}\sum_{j=1}^{c_{L-1}}{\left\|{\bm{w}}^{(L-1)}_{(j,:)}\right\|^{2}}\left(\sum_{i=1}^{m}{\epsilon}_{i}\sigma\left(\frac{1}{\left\|{\bm{w}}^{(L-1)}_{(j,:)}\right\|}\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(j,k)}^{(L-1),\top}\Pi_{g}{\bm{x}}^{(L-2)}_{i}{(k)}\right)\right)^{2}\right)$
$\leq\sum_{g=1}^{|G|}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\sum_{j=1}^{c_{L-1}}{\left\|{\bm{w}}^{(L-1)}_{(j,:)}\right\|^{2}}\left(\sum_{i=1}^{m}{\epsilon}_{i}\sigma\left(\frac{1}{\left\|{\bm{w}}^{(L-1)}_{(j,:)}\right\|}\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(j,k)}^{(L-1),\top}\Pi_{g}{\bm{x}}^{(L-2)}_{i}{(k)}\right)\right)^{2}\right)$
$\leq|G|M_{L-1}^{2}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-2]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}\left({\bm{x}}^{(L-2)}_{i}\right)\right\|^{2}}\right).$
Note that the last step involves the very same peeling argument for each term in the sum, and we removed it. Doing this iteratively, we can peel off all the layers, and get the final result. It is worth noting that the authors [golowich_size-independent_2018] provide a way of converting exponential depth dependence to polynomial dependence. Using this technique we can change the depth dependence from $|G|^{L-1/2}$ to $(L-1)\log|G|$ for $L>1$. See Section 3 of [golowich_size-independent_2018] for more discussions.
C.6.1 A generalization bound for gradual pooling
In this part, we consider a version of the multi-layer network with gradual pooling. We consider the following model:
$\hat{h}_{{\bm{u}},\{{\bm{w}}^{(l)},l\in[L]\}}({\bm{x}}):={\bm{u}}^{\top}P_{L}\circ{\sigma({\bm{W}}^{(L)}\sigma({\bm{W}}^{(L-1)}\dots P_{1}\circ\sigma({\bm{W}}^{(1)}){\bm{x}}\dots)},$ (29)
where the pooling operation changes the group size of each layer gradually to 1 as follows:
[…]
In other words, the pooling layer at layer $l$ pools from the cosets $G_{l}:=\hat{G}_{l-1}/\hat{G}_{l}$, for example by averaging out the elements of each left cosets. It can be seen as moving the operation from the group $\hat{G}_{l-1}$ to $\hat{G}_{l}$. This also means that:
$G=|\hat{G}_{0}|=|G_{1}||\hat{G}_{1}|=|G_{1}||G_{2}||\hat{G}_{2}|=\dots=|G_{1}|\times\dots\times|G_{L}|.$
Define the corresponding hypothesis space of functions with gradual pooling as follows:
${{\mathcal{H}}^{(L-g.p.)}:=\left\{\hat{h}_{{\bm{u}},\{{\bm{w}}^{(l)},l\in[L]\}}:\left\|{\bm{u}}\right\|\leq M_{1},\left\|{\bm{w}}_{i}\right\|\leq M_{i+1},i\in[L]\right\}.}$ (30)
We have the following theorem for this network.
Theorem C.9.
Consider the hypothesis space in eq. 30 consisting of functions defined in eq. 29. With probability at least $1-\delta$ and for all $h\in{\mathcal{H}}^{(L-g.p.)}$, we have:
${\mathcal{L}(h)\leq\hat{\mathcal{L}}(h)+2\left(\prod_{l=1}^{L}|\hat{G}_{l}|\right)\frac{b_{x}M_{1}M_{2}\dots M_{L+1}}{\sqrt{m}}+4\sqrt{\frac{2\log(4/\delta)}{m}}.}$
Proof.
For the rest of the proof, for simplicity, we assume that the elements of $G_{l}=\hat{G}_{l-1}/\hat{G}_{l}$ is represented by a member in $\hat{G}_{l-1}$ such that any element in $\hat{G}_{l-1}$ can be written uniquely as $g^{\prime}.g$ where $g^{\prime}=\hat{G}_{l},g\in{G}_{l}$.
The key for the proof is the intuitive observation that at each layer, we could use the group size utilized at that layer, more formally:
$\mathbb{E}_{\bm{\epsilon}}$ $\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}P_{L-1}\circ\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(1,k)}^{(L-1)}\circledast{\bm{x}}^{(L-2)}_{i}{(k)}\right)\right\|^{2}\right)$
$=\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\sum_{g^{\prime}\in\hat{G_{L-1}}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\frac{1}{G_{L-1}}\sum_{g=1}^{G_{L-1}}\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(1,k)}^{(L-1),\top}\Pi^{L-1}_{g^{\prime}g}{\bm{x}}^{(L-2)}_{i}{(k)}\right)\right\|^{2}\right)$
$\leq\frac{1}{G_{L-1}}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\sum_{g^{\prime}\in\hat{G_{L-1}}}\sum_{g=1}^{G_{L-1}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(1,k)}^{(L-1),\top}\Pi^{L-1}_{g^{\prime}g}{\bm{x}}^{(L-2)}_{i}{(k)}\right)\right\|^{2}\right)$
$\leq\frac{1}{G_{L-1}}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\sum_{g\in\hat{G_{L-2}}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(1,k)}^{(L-1),\top}\Pi^{L-1}_{g}{\bm{x}}^{(L-2)}_{i}{(k)}\right)\right\|^{2}\right)$
$\leq\frac{\hat{G}_{L-2}}{G_{L-1}}\left\|{\bm{w}}_{(1,:)}^{(L-1)}\right\|^{2}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-2]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}\left({\bm{x}}^{(L-2)}_{i}\right)\right\|^{2}}\right)$
$\leq\hat{G}_{L-1}\left\|{\bm{w}}_{(1,:)}^{(L-1)}\right\|^{2}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-2]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}\left({\bm{x}}^{(L-2)}_{i}\right)\right\|^{2}}\right).$
Using the above argument, we can revisit the proof above and see that:
$\mathbb{E}_{\bm{\epsilon}}$ $\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}{\bm{x}}^{(L-1)}_{i}\right\|^{2}}\right)$
$=\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-1]\}}\left\|\sum_{i=1}^{m}{\epsilon}_{i}\begin{pmatrix}P_{L-1}\circ\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(1,k)}^{(L-1),\top}\circledast{\bm{x}}^{(L-2)}_{i}{(k)}\right)\\ \vdots\\ P_{L-1}\circ\sigma\left(\sum_{k=1}^{c_{L-2}}{\bm{w}}_{(c_{L-1},k)}^{(L-1),\top}\circledast{\bm{x}}^{(L-2)}_{i}{(k)}\right)\\ \end{pmatrix}\right\|^{2}\right)$
$\leq|\hat{G}_{L-1}|M_{L-1}^{2}\mathbb{E}_{\bm{\epsilon}}\left(\sup_{\{{\bm{w}}^{(l)},l\in[L-2]\}}{\left\|\sum_{i=1}^{m}{\epsilon}_{i}\left({\bm{x}}^{(L-2)}_{i}\right)\right\|^{2}}\right).$
We omitted many steps in the proof as they are exactly similar to the proof above. ∎
Remark C.10.
We would like to highlight that the gradual pooling breaks the equivariance with respect to the original $G$, and it is not a common practice in geometric deep learning to do it, in contrast with typical convolutional neural networks.
@@ Appendix F Frequency Domain Analysis
For the hypothesis space ${\mathcal{H}}$ defined in eq. 4, the average pooling operation, $\sigma(\cdot)$ as a $1$-Lipschitz positively homogeneous activation function, the generalization error is bounded as ${\mathcal{O}}(\frac{b_{x}M_{1}M_{2}}{\sqrt{m}})$.
@@
where we used the positively homogeneity property of the activation function, and we defined:
@@ Appendix G Proofs for Weight Sharing
The following works do not consider the generalization error for equivariant networks. However, the convolutional networks, which are studied in their work, are closely related to our results. We have also summarized our results and their comparison in Table 2.
Reference Bound
Equivariance (Theorem 4.1) $\frac{b_{x}M_{1}M_{2}}{\sqrt{m}}$
Weight sharing (Proposition 6.1) $\frac{b_{x}M_{1}M^{w.s.}_{2}}{\sqrt{m}}$
Locality (Proposition 7.1) $\sqrt{\frac{O_{\Phi}}{|G|}}\frac{b_{x}M_{1}M_{2}}{\sqrt{m}}$
Non-Equivariant Networks [golowich_size-independent_2018] $\frac{b_{x}M_{1}\left\|{\bm{W}}\right\|_{F}}{\sqrt{m}}$
Weight sharing, orthogonal ${\bm{b}}_{k,l},k\in[|G|]$ $\frac{b_{x}M_{1}M_{2}}{\sqrt{m}}$
$B$-sparse filters (Proposition 7.2 $b_{x}M_{1}M_{2}/\sqrt{mB}$
Linear convolution [vardi_sample_2022] $\sqrt{O_{\Phi}}\frac{b_{x}M_{1}\left\|{\bm{W}}\right\|_{2\to 2}}{\sqrt{m}}$
CNN with Pooling [vardi_sample_2022] $\frac{b_{x}\left\|{\bm{W}}\right\|_{2\to 2}{\color[rgb]{0,0,1}\log m\sqrt{\log(mc_{1}|G|)}}}{\sqrt{m}}$
[graf_measuring_2022] ${\mathcal{O}}\left(\frac{b_{x}M_{2}\log(|G|{\color[rgb]{0,0,1}\max(c_{0},c_{1}))}\sqrt{\left\|{\bm{W}}\right\|_{2\to 2}+\left\|{\bm{w}}\right\|_{2,1}}}{\sqrt{m}}\right)$
Table 2: The table summarizes the main results of the paper and compares with the selected prior works. The top three rows are the main results of the paper expressed in their full generality (Reminder : $M^{w.s.}_{2}=\max_{l\in[|G|]}\left\|\left(\sum_{k=1}^{|G|}{\bm{w}}_{(j,c)}(k){\bm{b}}_{k,l}^{\top}\right)_{c\in[c_{0}],j\in[c_{1}]}\right\|_{F}$). Next three rows summarize the result for the case where the network is not equivariant, the optimal weight sharing and the band-limited filters. The bound from the most relevant results are also summarized in the last rows of the table. The explicit dimension dependency of these bounds are highlighted with a different color.
@@
Our work is closely related to [vardi_sample_2022]. We focus on equivariant networks and upper bounds on the sample complexity. The networks considered in this work are slightly more general with per-channel linear and non-linear pooling operations. They also provide a bound in Theorem 7 of their paper for pooling operations $\rho:\mathbb{R}^{|G|}\to\mathbb{R}$ that are 1-Lipschitz with respect to the $\ell_{\infty}$ norm. This class includes average and max pooling operations. Their bound has logarithmic dimension dependence, the artifact of using the covering number-based arguments (see below for more discussions). The bound is also dependent on the spectral norm of the matrix ${\bm{W}}$ instead of the norm of the weight ${\bm{w}}$. It is not clear if the dependence on the spectral norm can be removed in their proof. For local filters, they provide a dimension-free bound that also depends on the spectral norm of ${\bm{W}}$. However, this time, the dependence on the spectral norm can be substituted with the norm of the weight vector, as it is evident from their proof. In general, all our bounds depend only on the norms of the weight vector, which is tighter than using Frobenius or spectral norm of ${\bm{W}}$. The authors in [vardi_sample_2022] provide lower bounds on the sample complexity, which we relegate to future works.
@@
The paper of [graf_measuring_2022] relies mainly on Maurey’s empirical lemma and covers a number of arguments. Rademacher complexity bounds either work directly on bounding the Rademacher sum through a set of inequalities or utilize chaining arguments and Dudley’s integral inequality. Dudley’s integral requires an estimate of the covering number. The main techniques of finding the covering number, like volumetric, Sudakov, and Maurey’s lemma techniques, manifest different dimension dependencies, with the latter usually providing the mildest dependence. Completely dimension-free bounds are obtained mainly via the direct analysis of the Rademacher sum. Therefore, the bounds in [graf_measuring_2022] are a mixture of dimension-dependent and norm-dependent terms. Their bounds are, however, quite general and consider different activation functions and residual connections.
@@
In [wang_theoretical_2023], they study the same set of problems as ours, related to locality and weight sharing, but from different perspectives. First, the approximation theory perspective, studying the class of functions a model can approximate, is complementary to our paper on the generalization error using statistical learning theory. The learning theory part of the paper is focused on a particular "separation task" (see (5)) with a fixed input distribution. They also use Rademacher complexity but bound it with a covering number. We work with general input distributions and arbitrary tasks. Besides, our bounds are dimension independent, while their bound explicitly depends on the input dimension, an artifact of using covering number for bounding RC.
@@
In [li_why_2021], the authors consider the sample complexity of convolutional and fully connected models and construct a distribution for which there is a fundamental gap between them in the sample complexity. Their analysis is based on the VC dimension and uses Benedek-Itai’s lower bound. Their final bound is, therefore, dependent on the input dimension. In contrast, one of the main interests of our paper was to explore dimension-free bounds. We additionally study the impact of locality and weight sharing.
@@
Dimension-free bounds are interesting because they hint at why the generalization error is unaffected by over-parametrizing the model. In this sense, the notion of dimension refers mainly to the input dimension, the number of channels, the width of a layer, and the number of layers. The same motivation existed for obtaining norm-based bounds for RKHS SVM models. The works of [golowich_size-independent_2018] and [vardi_sample_2022] are two recent examples, with the former work providing an extensive review of other dimension-free bounds. Choosing a proper norm to get dimension-free bounds is one of the main questions in learning theory. Naturally, this is easier to get for larger norms like Frobenius norms, and the question is if we can have similar dimension-free results for smaller norms like spectral or, as we show in the paper, the norm of effective parameters.
@@
In the literature around deep learning theory, there are many works questioning the relevance of classical learning theoretic results, for example, using Rademacher complexity, norm-based bounds, and uniform convergence. We can refer, for example, to works such as [nagarajan_uniform_2019, jiang_fantastic_2020, zhang_understanding_2017, zhu_understanding_2021]. It is fair to ask if Rademacher complexity analysis, in light of these criticisms, can provide tight bounds for our case. First of all, the preceding works focused on state-of-the-art deep neural networks with a large number of layers. Our study is limited to single-hidden layer neural networks. Second, Rademacher complexity is indeed tight for support vector machines [Steinwart2008-jo, shalev-shwartz_understanding_2014, Bartlett2002-gt]. If we fix the first layer of our network and only train the last layer, then we have a linear model, and it is known that training it with Gradient descent will lead to minimum $\ell_{2}$-norm solutions. For this sub-class of learning algorithms, the analysis boils to previous works on minimum norm classifiers as in [shalev-shwartz_understanding_2014] for which the Rademacher complexity analysis is tight. Therefore, we believe that Rademacher complexity can still be a relevant tool for shallower models, including those discussed in this paper.
@@
Previous works also tried to characterize the effect of equivairance on the training dynamics. [lawrence2022implicit] showed that deep equivariant linear networks trained with gradient descent provably converge to solutions that are sparse in the Fourier domain. While this work focuses on the effect of equivariance on optimization, the authors suggest that this implicit bias towards sparse Fourier solutions can be beneficial towards generalization in bandlimited datasets - which is common for non-discrete compact groups such as rotation groups and a common assumption in steerable CNNs. Similar results were also previously obtained in [gunasekar2018implicit] for standard convolutional networks, which can be seen as an instance of GCNN with discrete abelian groups.
@@
This paper provided a Rademacher complexity-based bound for single-hidden layer equivariant networks. We obtained a lower bound, which showed the tightness of our analysis. The numerical results showed compelling scaling behaviors for our bound. However, there is still a gap between our bound and the generalization error. We conjecture that the gap is related to the idea of implicit bias. We conjecture that stochastic gradient descent training implicitly biases toward only a subset of the hypothesis space even smaller than norm-bounded functions. It is an important next step to tie these two analyses together by characterizing the implicit bias, for example, in terms of some kind of norms, and then using Rademacher complexity analysis to find a bound on the generalization error based on the implicit bias.
On the other hand, we provided a lower bound on the Rademacher complexity. It would be interesting to obtain such bounds by finding the fat-shattering dimension. If two points in the dataset can be transformed into each other by the action of group $G$, then the space of $G$-invariant functions cannot shatter such datasets. Therefore, the data points should be picked on different orbits so they can be shattered. Such construction will be interesting in the future.